which of the following units belongs to magnetic flux

a. Amper volt
b. Tesla
c. Volt / s
d. Tesla m
e. C. Ω
Answer #1

Magnetic flux is defined as no of magnetic field lines through a surface.

Magnetic flux=Magnetic*area

unit of magnetic flux = unit of magnetic field * unit of area

\begin{align*} &=\left (\text{unit of }\frac{Force}{current * length} \right )* m^{2}\\ &=\frac{N}{A*m}*m^{2}\\ &=\frac{N*m}{A}\\ &=\frac{N*m*s}{C}\text{ [since, ampere = C/s]}\\ &=V.s \text{ [since volt = J/C = Nm/C]} \end{align*}

Therefore, Unit of magnetic flux in terms of

a) tesla is T.m^{2}

b) volt is \text{V.s or volt.s}

c) joule is \frac{J}{A} \text{ [where A = ampere}]

so, none of the options is matching.